OpenStudy (anonymous):

write the slope intercept equation for the line that passes through. (-7,-1) and has a slope -6/7 plezz show work

OpenStudy (dumbcow):

y = mx+b my x-value is -7 my y-value is -1 -1 = (-6/7)*-7 +b -1 = 6 +b -7 = b y = (-6/7)x - 7